Tickets to a local movie Were sold at $6 for adults and $4.50 for students. if 270 tickets were sold for a total of $1395. how many tickets were sold of each kind?
Answer by meay7c(10)

For this problem we will identify the number of adult and student tickets.
We will assign the following variables:
Let x = number of adult tickets sold.
Let y = number of student tickets sold.
Here we can develop the following equations:
6x + 4.50y = 1395 (this equations tells us the amount of sales in dollars)
x + y = 270 (this equation tells us the total units sold)
We can apply the elimination method used in a system of equations.
6x + 4.50y = 1395 (Equation 1)
x + y = 270 (Equation 2)
Lets multiply equation 2 by (-6) to eliminate x, thus solving for y.
We arrive at:
6x + 4.50y = 1395
-6x - 6y = -1620
___________________
-1.50 y = -225
Divide both sides by -1.50. Yielding:
y = 150
Next take y(150) and plug into equation 2 to solve for x.
x + 150 = 270
x = 120
Check your solutions:
120 + 150 = 270 Yes!!